# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ 5.2 Examples: Self-Normalization, Bernstein, and Hoeffding
Under the same assumptions as Example 4.8, for any stopping time $\tau$, $\alpha\in(0,1)$, and a super-uniform $U$ independent from $\mathcal{F}_{\infty}$,
$\mathbb{P}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{\tau}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{\tau}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\geqslant dU/\alpha\right)\leqslant\alpha.$ (73)
The bound above, as those matrix supermartingales in Section 4.2, does not bear closed-form confidence sets but still implies a computable randomized sequential test for the null, e.g., that the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})$ all equal $\mathbf{M}$ with marginal type I error rate within $\alpha$.
However, a closed-form concentration can be computed when further assumption is introduced on top of the finite variance in Example 4.8. This is the Hoeffding-type e-process by Howard et al. [2020, Lemma 3], which we present as a Bernstein-type result akin to Mackey et al. [2014, Corollary 4.2] for generality.
Example 5.4 (Matrix Bernstein Scalar e-Process).
Under the same assumptions as Example 4.8, suppose further that $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ almost surely for some predictable sequence $\{\mathbf{B}_{n}\}$. Then, the self-normalized supermartingale $\{L_{n}^{\textsf{SN}}\}$ upper bounds the following process:
$L_{n}^{\textsf{SN}}\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\}.$ (74)
@@
Taking $\psi(x)=x^{2}/2$, $\mathbf{C}_{n}=\frac{1}{3}(\mathbf{X}_{n}-\mathbf{M})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{2}{3}\mathbf{V}_{n}$ in Lemma 5.1, from (70),
$L_{n}^{\textsf{SN}}$ $\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\right\}$ (75)
$\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\},$ (76)
concluding the proof. ∎
A closed-form, randomized, stopping time-valid matrix Bernstein or Hoeffding inequality on the concentration of weighted sample average is immediately in force by applying (UVI) to $\{L_{n}^{\textsf{SN}}\}$. Below, we denote by $\overline{\mathbf{X}}_{n}^{\gamma}$ the weighted average $\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{n}\mathbf{X}_{n}}{\gamma_{1}+\dots+\gamma_{n}}$ w.r.t. some positive weight sequence $\{\gamma_{n}\}$.
Theorem 5.5 (Uniformly Randomized Stopped Matrix Bernstein/Hoeffding Inequality).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$, $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ and $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ almost surely for predictable $\{\mathbf{B}_{n}\}$ and $\{\mathbf{V}_{n}\}$. Then, for any stopping time $\tau$, predictable positive sequence $\{\gamma_{n}\}$, $\alpha\in(0,1)$, and super-uniform random scalar $U$ independent of $\mathcal{F}_{\infty}$,
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{6}\lambda_{\max}\left(\sum{\gamma_{i}^{2}}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$$\leqslant$ (77)
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}\left(\sum_{i=1}^{\tau}\gamma_{i}^{2}\mathbf{B}_{i}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$ $\leqslant\alpha.$ (78)
We now remark at length on this theorem. First, the Hoeffding term $\frac{1}{2}\lambda_{\max}\left(\sum\gamma_{i}^{2}\mathbf{B}_{i}\right)$ comes from applying $\mathbf{V}_{i}\preceq\mathbf{B}_{i}$ to the Bernstein term $\frac{1}{6}\lambda_{\max}\left(\sum\gamma_{i}^{2}(\mathbf{B}_{i}^{2}+2\mathbf{V}_{i})\right)$, among which one may choose, as is with scalar Bernstein versus Hoeffding, depending on whether variance bounds $\{\mathbf{V}_{n}\}$ better than the total square bounds $\{\mathbf{B}_{n}\}$ are known. Second, the strength of this bound can be seen from the following fact: taking $\{\mathbf{M}_{n}\}$ to be a constant $\mathbf{M}$, $\{\mathbf{B}_{n}\}$ a constant $\mathbf{B}$, $\{\mathbf{V}_{n}\}$ a constant $\mathbf{V}$, $\tau$ a deterministic sample size $n$ and letting $U=1$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$, (77) reads
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B}+2\mathbf{V})}{3n}}\right)\leqslant\alpha,$ (79)
where $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. This even improves the optimal matrix Hoeffding inequality by Mackey et al. [2014, Corollary 4.2] whose dependence on $\mathbf{B}$ and $\mathbf{V}$ is $\frac{1}{2}\lambda_{\max}(\mathbf{B}+\mathbf{V})$ compared to our smaller $\frac{1}{3}\lambda_{\max}(\mathbf{B}+2\mathbf{V})$, [^4] The optimality of Mackey et al. [2014, Corollary 4.2] is in terms of the overall multiplicative constant which (79) achieves as well, not the distribution of the constant between $\mathbf{B}$ and $\mathbf{V}$. Even tighter dependence in this regard shall be presented soon in Section 5.3. and recovers (MHI), both via $\mathbf{V}\preceq\mathbf{B}$. The uncompromised inequality (77) further allows for stopping time-validity, time-varying and adaptive weighting, and randomized tightening simultaneously. Third, in the i.i.d. (or homoscedastic more generally) sequential (sample size $\tau$ not fixed) case, following Waudby-Smith and Ramdas [2020], Wang and Ramdas [2023a], the weights $\{\gamma_{n}\}$ can be set as a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$. By doing so, the concentration tightens itself as time goes on; or, the corresponding sequential test has asymptotic power 1 [Waudby-Smith and Ramdas, 2023, Table 1]. Moreover, it can be seen that the closed-form concentration of $\overline{\mathbf{X}}_{\tau}$ here is achieved via multiple steps of scalarization, each step loosening the bound by a little. Therefore, if one does not want to quantify uncertainty but is only interested in sequentially testing if the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$, the more powerful test in the current subsection is still with the largest process, the self-normalized supermartingale $\{L^{\textsf{SN}}_{n}\}$ (i.e., rejecting whenever $L^{\textsf{SN}}_{n}$ surpasses $Ud/\alpha$), with variance upper bounds $\mathbf{V}_{n}$ supplied by the upper bounds $\mathbf{B}_{n}$.
5.3 Matrix Empirical Bernstein Inequality
We now present a closed-form empirical Bernstein inequality that enjoys the same stopping time-validity, adaptive weighting, and randomization as well, foreshadowed by a matrix expectation bound in Proposition 4.10. Empirical Bernstein inequalities for scalars differ from “classical” Bernstein inequalities, e.g. our (77) when $d=1$, in that the population variance is replaced by an empirical variance in the deviation bound. They therefore achieve the advantage of Bernstein’s inequalities over Hoeffding’s inequalities without requiring knowing the variance beforehand.
Scalar empirical Bernstein inequalities are often derived in two different ways. First, a union bound between a classical Bernstein inequality and a concentration inequality on the empirical variance, a method that early empirical Bernstein results [Audibert et al., 2009, Maurer and Pontil, 2009] all share. Second, a self-normalization technique pioneered by Howard et al. [2021, Theorem 4] and later followed by Waudby-Smith and Ramdas [2023], Chugg et al. [2023], among others, which enables supermartingale-based, stopping time-valid concentration, and variance proxy by predictable estimates other than the empirical variance. These two methods are inherently different and as argued convincingly by Howard et al. [2021, Appendix A.8]: the latter’s avoidance of the union bound produces a sharper concentration. No matrix empirical Bernstein results, especially those in the second fashion, are previously known.
Our matrix extension of Howard et al. [2021, Theorem 4] starts from the bound we realier prove as Proposition 4.10. Recall that $\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ for $x\in[0,1)$, and that
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n}))$ (80)
under, among other, a key assumption that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$. We can use Lemma 5.1 again to produce a scalar supermartingale and a closed-form spectral concentration inequality.
Theorem 5.6 (Uniformly Randomized Stopped Matrix Empirical Bernstein Inequality).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$. Let $\{\widehat{\mathbf{X}}_{n}\}$ be a sequence of predictable and integrable $\mathcal{S}_{d}$-valued random matrices such that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ almost surely. Then, for any predictable $(0,1)$-valued sequence $\{\gamma_{n}\}$,
$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)$ (81)
is a supermartingale. Further, for any stopping time $\tau$, $\alpha\in(0,1)$, and super-uniform scalar random variable $U$ independent of $\mathcal{F}_{\infty}$, we have
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha.$ (82)
Before we remark on the uncompromised Theorem 5.6, let us first write down its fixed-time, non-randomized, fine-tuned special case with $\tau=n$ and $U=1$ which shall justify the “empirical Bernstein” name it bears.
Corollary 5.7 (Matrix Empirical Bernstein Inequality).
Suppose $\alpha\in(0,1)$. Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be $\mathcal{S}_{d}^{[0,1]}$-valued i.i.d. random matrices with mean $\mathbf{M}=\mathbb{E}\mathbf{X}_{1}$. Let $\overline{\mathbf{X}}_{i}=\frac{1}{i}(\mathbf{X}_{1}+\dots+\mathbf{X}_{i})$ and $\overline{\mathbf{X}}_{0}=0$. Then, for any predictable $(0,1)$-valued sequence $\gamma_{1},\dots,\gamma_{n}$,
[…]
Further, if we define the following variance proxies
$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (84)
$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right),$ (85)
and set $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$ for $i=1,\dots,n$, then,
$D_{n}^{\mathsf{EB}}\leqslant\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}\quad\approx\sqrt{\frac{9\log(d/\alpha)\lambda_{\max}(\mathbf{V})}{2n}}$ (86)
where $\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1})$; further,
$\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}\quad\text{almost surely}.$ (87)
We prove both Theorem 5.6 and Corollary 5.7 in Section B.4. The asymptotic behavior (87) of deviation bound $D_{n}^{\mathsf{EB}}$ is striking as it adapts fully to, without knowing, the true variance $\mathbf{V}$, a strict improvement even compared to the deviation bound $D_{n}^{\mathsf{B}}$ of the our previous matrix Bernstein inequality (79),
$\sqrt{n}D_{n}^{\mathsf{B}}=\sqrt{2\log(d/\alpha){\frac{1+2\lambda_{\max}(\mathbf{V})}{3}}},$ (88)
which requires knowing beforehand, but only adapts partially to, the variance $\mathbf{V}$. Indeed, when $\mathbf{V}$ is small, (88) can be arbitrarily larger than (87). We can also compare Corollary 5.7 to another matrix Bernstein inequality in the literature by Tropp [2012, Theorem 1.4], who proves that $\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant D_{n}^{\mathsf{TB}}\right)\leqslant\alpha$ under the same assumptions (plus knowing the variance $\mathbf{V}$), where its deviation bound
$D_{n}^{\mathsf{TB}}=\frac{\log(d/\alpha)+\sqrt{\log^{2}(d/\alpha)+18\log(d/\alpha)n\lambda_{\max}(\mathbf{V})}}{3n}$ (89)
satisfies the same asymptotic $\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{TB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$ as (87), which we achieve without requiring knowing $\mathbf{V}$. Other matrix Bernstein bounds in the literature, like the ones by Mackey et al. [2014, Corollary 5.2] and by Oliveira [2009], are looser than the Tropp [2012, Theorem 1.4] bound we discuss here as acknowledged by Mackey et al. [2014] themselves.
Our Theorem 5.6 and Corollary 5.7 owe much to the techniques developed by Waudby-Smith and Ramdas [2023, Theorem 2 and Remark 1] in the scalar case (who in turn build on the earlier result by Howard et al. [2021, Theorem 4] via the “predictable mixing” sequence $\{\gamma_{n}\}$). We recover exactly as special cases those two known results when $d=1$.
We reiterate our remark following Proposition 4.10 that $\widehat{\mathbf{X}}_{n}$ in Theorem 5.6 can usually be taken as the empirical average from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, which is exactly what is done in Corollary 5.7. In the i.i.d. (or homoscedastic more generally) sequential case without the sample size $n$ fixed in advance, we can again (recalling a similar discussion following Theorem 5.5) set the weights $\{\gamma_{n}\}$ to be a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{n-1}}}$. Under this weight sequence, we see the choice of a weighted average $\widehat{\mathbf{X}}_{n}=\overline{\mathbf{X}}_{n-1}^{\psi_{E}(\gamma)}$ is more reasonable as it roughly minimizes the weighted sum of squares $\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}$ in (82) into a weighted empirical variance — a heuristic which should work for Proposition 4.10 as well. Obviously, as long as $\widehat{\mathbf{X}}_{n}$ is any average, weighted or not, of $\mathbf{X}_{1},\dots,\mathbf{X}_{n-1}$, the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is met when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$.
@@ 6 Experiments: To Scalarize or not to Scalarize
We have developed thus far two categories of matrix concentration bounds: those obtained via the pure matrix approach, typically expressed in the $\preceq$ order, in Sections 3 and 4; and those obtained via scalarization, i.e. taking the trace, in Section 5. It is therefore a natural question to ask if either of these is tighter than the other.
A fair comparison can be raised between the self-normalized matrix and scalar test supermartingales, Example 4.8 and Example 5.2 — for both are under the assumption of finite variance and based on the key one-step self-normalized bound (71). We place them in the setting of the following sequential testing problem: let $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ be an i.i.d. sequence of matrices in $\mathcal{S}_{d}$, with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$. We would like to sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$. Let us call the tests derived from Example 4.8 and Example 5.2 for this null Matrix and Scalar respectively. Ignoring the randomizer for simplicity (plug in $\mathbf{U}=\mathbf{I}$), Matrix computes the matrix-valued process
@@
rejecting when $\mathbf{Y}_{n}^{\textsf{SN}}\npreceq\mathbf{A}$ for some threshold matrix $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$; Scalar computes the scalar-valued process

### v4

@@ 5.3 Randomized Ahlswede-Winter Bounds
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Let $U$ be a super-uniform random scalar independent from $\mathcal{F}_{\infty}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1});$ (75)
and
$\mathbb{P}(\exists n<\tau,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\text{ or }\mathbf{Y}_{\tau}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (76)
@@
First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Theorem 3.3. We apply Theorem 5.3 to obtain the first inequality. For the second inequality, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$ and apply (75) to the stopping time $\nu\wedge\tau$. ∎
These two inequalities above, we note, generalize respectively the stopped and time-uniform statements of (MVI), just as the two inequalities by Ramdas and Manole [2024, Theorem 4.1, Corollary 4.1.1] generalize the two forms of the scalar (VI). It is important to note that while it is tempting to hope for the following randomized time-uniform (MVI):
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}),$ (77)
the temptation must be resisted as it is known to fail in the scalar case; see the discussion before Corollary 4.1.1 by Ramdas and Manole [2024]. Theorem 3.4 provides the following improvement over Algorithm 1 where the user may prespecify a stopping criterion $\tau$ beforehand. As is already stated in Algorithm 1, whenever $\mathbf{L}\mathbf{R}\npreceq\mathbf{A}$, $\mathcal{H}_{0}$ is rejected; what is now different is that when the stopping criterion $\tau$ is met and $\mathcal{H}_{0}$ is not yet rejected, the user can draw a $U\sim\operatorname{Unif}_{(0,1)}$ to see if $\mathbf{L}\mathbf{R}\npreceq U\mathbf{A}$, and reject $\mathcal{H}_{0}$ when this happens.
@@ 6 Experiments: To Scalarize or Not to Scalarize
We now return to our discussion in Section 5.2 that from the condition
$\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime})$ (78)
that holds under the null, one can either use our approach to construct a matrix test supermartingale via Lemma 3.5, or use the stepwise scalarization technique formalized as Lemma 5.2 to construct a scalar test supermartingale. We now compare these two empirically.
We conduct the comparison with the self-normalized Example 3.8, and consider the two corresponding sequential test algorithms, Matrix and Scalar. More specifically, both receive the stream of i.i.d matrices $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ in $\mathcal{S}_{d}$ with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$ and sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$. Matrix computes the matrix-valued process
@@
rejecting when $\mathbf{Y}_{n}^{\textsf{SN}}\npreceq\mathbf{A}$ for some threshold matrix $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$ picked beforehand; Scalar computes the scalar-valued process